# Distributed Harmonization: Federated Clustered Batch Effect Adjustment and Generalization

### v1

@@ Abstract.
Independent and identically distributed (i.i.d.) data is essential to many data analysis and modeling techniques. In the medical domain, collecting data from multiple sites or institutions is a common strategy that guarantees sufficient clinical diversity, determined by the decentralized nature of medical data. However, data from various sites are easily biased by the local environment or facilities, thereby violating the i.i.d. rule. A common strategy is to harmonize the site bias while retaining important biological information. The ComBat is among the most popular harmonization approaches and has recently been extended to handle distributed sites. However, when faced with situations involving newly joined sites in training or evaluating data from unknown/unseen sites, ComBat lacks compatibility and requires retraining with data from all the sites. The retraining leads to significant computational and logistic overhead that is usually prohibitive. In this work, we develop a novel Cluster ComBat harmonization algorithm, which leverages cluster patterns of the data in different sites and greatly advances the usability of ComBat harmonization. We use extensive simulation and real medical imaging data from ADNI to demonstrate the superiority of the proposed approach.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Imagingccs: Computing methodologies Distributed algorithmsccs: Computing methodologies Machine learning
@@ 1. Introduction
The recent advances in machine learning approaches have greatly advanced biomedical data analysis. In the brain imaging analysis, for example, Magnetic Resonance Imaging (MRI) has been used for the detection and disease progression of many diseases, such as Mild Cognitive Impairment (MCI) (Wang et al., 2018), Parkinson’s disease (Garcia Santa Cruz et al., 2023), and Brain Tumor Detection (Abdusalomov et al., 2023). However, one critical challenge with brain imaging is that the brain imaging is sensitive to scanner or protocol effect (Zhan et al., 2014; Pfaehler et al., 2019), also commonly referred to as site effect or batch effect, leading to the fact that brain imaging from multiple sites is not independent and identically distributed (i.i.d.). The bias in non-i.i.d. data will cause unstable prediction performance and poor generalization performance to unseen data (Yang et al., 2021). Consequently, developing an algorithm that can eliminate these types of bias ensures consistent and reliable outcomes in the deployment of machine learning models within the medical imaging domain.
@@
ComBat (Johnson et al., 2006) is a well-known harmonization technique and has been shown to be helpful in mitigating the site effect of neuroimaging data introduced by multiple sites sampling (Orlhac et al., 2021). Despite its utility, one of the central ideas is that ComBat debias the site effect independently according to the (local) site data, which induces one critical limitation, its inability to evaluate site effects coming from unseen or unknown sites without undergoing a retraining process. The requirement of retraining is hindered by substantial computational costs when comes to real-world deployment, especially when dealing with large datasets from multiple sites. This limitation underscores the need for a more efficient and broadly applicable approach in mitigating site effects of medical data, especially neuroimaging data.
@@
Furthermore, a centralized setting for site effect harmonization introduces extra concerns. For instance, sharing data directly among multiple sites to apply ComBat harmonization poses challenges to the security of confidential data and the protection of patient privacy. So direct training on all the data is often impractical in the medical domain. This underscores the need to develop harmonization algorithms in a decentralized manner that can still effectively harmonize without gathering data from all sites while still maintaining competitive performance in the centralized setting. Particularly, the Distributed ComBat (Chen et al., 2022), a distributed version of the ComBat harmonization algorithm, has demonstrated its harmonization capability meanwhile obeying decentralized manners. Nevertheless, Distributed ComBat suffers from the same limitation as the original ComBat, i.e., it cannot estimate site effects from unseen sites without retraining.
@@
Because a significant part of site effects in medical data are ultimately rooted in medical instruments, e.g., MRI scanners from different manufacturers and configurations, the bias underneath the sites may not be independent and may exist clustering structures. In this paper, we proposed the Cluster ComBat method, an extension of the original ComBat algorithms that leverages the cluster patterns of site effect. This approach enables the estimation of site effects from unknown sites without necessitating for retraining process. Furthermore, we also developed a distributed version Cluster ComBat that demonstrates its efficacy in harmonizing data obeying decentralized manners. Our empirical findings show that Cluster ComBat in both centralized and decentralized settings perform better than their respective counterparts under both synthetic and real-world neuroimaging datasets.
@@ 2. Related Works
Distributed Learning. Preserving the privacy of users’ information is a crucial issue that needs to be considered as an important aspect to evaluate in a learning algorithm  (Li et al., 2021; Fukasawa et al., 2004). Especially in a healthcare setting, where the health data is sensitive, we need to design a distributed learning approach that avoids leaking any private information from hospitals’ data  (Welten et al., 2022; Wirth et al., 2021; Haque et al., 2023). For example, for brain imaging, federated learning, a distributed machine learning algorithm, has proven to be effective in analyzing neuroimaging for cognitive detection tasks while still protecting patients’ information  (Sandhu et al., 2023; Dipro et al., 2022). Moreover, for health records data, to avoid sharing raw data between sites, the use of first-order and second-order gradients of the likelihood function of sites has been shown to be sufficient for achieving high accuracy in classification task  (Duan et al., 2019). In addition, the distributed version of generalized linear mixed models (GLMM) can also achieve nearly identical results in analyzing electronic health records data as in a centralized setting  (Yan et al., 2023). Relating to harmonization methods in a decentralized setting, Distributed ComBat (Chen et al., 2022) and Federated Learning ComBat (Silva et al., 2023) have been developed to harmonize neuroimaging without the need for sharing information between hospitals.
@@ 3.1. Preliminary: ComBat Harmonization
(3) $\gamma_{ig}^{*}=\frac{N_{i}\overline{\tau}_{i}^{2}\hat{\gamma}_{ig}+\delta_{ig}^{2}\overline{\gamma}_{i}}{N_{i}\overline{\tau}_{i}^{2}+{\delta_{ig}^{*}}^{2}}\quad\quad\quad{\delta_{ig}^{*}}^{2}=\frac{\overline{\theta}_{i}+\frac{1}{2}\sum_{j}(Z_{ijg}-\gamma_{ig}^{*})^{2}}{\frac{1}{2}N_{i}+\overline{\lambda}_{i}-1},$
where $\overline{\tau}_{i}^{2}$, $\overline{\gamma}_{i}$, $\overline{\theta_{i}}$, $\overline{\lambda_{i}}$ are computed through method of moments.
@@ 3.2. Cluster ComBat
Using cluster-wise shared harmonization parameters, we can generalize knowledge from previous $N$ sites to the new unseen site once we know the cluster each data point from this site belongs to. Additionally, the estimation process of harmonization parameters $\gamma_{cg}$ and $\delta_{cg}$ can benefit from multiple sites’s data points within the same cluster, considering the sample number for estimating each parameter group is enlarged. And we name this algorithm Cluster ComBat.
@@
$\hat{\alpha}_{g},\hat{\beta}_{g},\hat{\delta}_{cg}^{*},\hat{\gamma}_{cg}^{*}$ harmonization parameters and $k$ trained K-means model
@@
$y_{ijg}$ unseen site’s data, $X_{ij}$ biological covariate of new client, previously estimated parameters $\hat{\alpha}_{g},\hat{\beta}_{g},\hat{\delta}_{cg}^{*},\hat{\gamma}_{cg}^{*}$, and trained K-means $k$
@@ 3.3. Distributed Cluster ComBat
The Distributed Cluster ComBat has the following steps:  i) each site estimates feature-wise parameters $\hat{\alpha}_{ig}$, $\hat{\beta}_{ig}$, and $\hat{\gamma}_{ig}$ locally at the same time, and sends parameters to the central server;  ii) the central server performs K-means clustering based on $\hat{\alpha}_{ig}$, $\hat{\beta}_{ig}$, and $\hat{\gamma}_{ig}$ for $i\in[M]$;  iii) the central server aggregates $\{\hat{\alpha}_{ig}\}_{i\in M}$, $\{\hat{\beta}_{ig}\}_{i\in M}$ and $\{\hat{\gamma}_{ig}\}_{i\in M}$ to estimates the global feature-wise parameters $\hat{\alpha}_{g}$, $\hat{\beta}_{g}$ and $\hat{\gamma}_{ig}$, and then sends back to each site;  iv) each site standardized the local data using global feature-wise parameters, then locally estimates $\hat{\delta}_{ig}$ and $\hat{\gamma}_{ig}$;  v) each site sends locally estimated harmonization parameters to the server;  vi) server aggregates harmonization parameters within each cluster to estimate the cluster-wise ones, then sends back to each site; the aggregation procedure precisely follow the procedure outlined in Figure 1 of the original Distributed ComBat algorithm paper (Chen et al., 2022);  vii) each site performs local harmonization based on cluster-wise harmonization parameters. The procedure is summarized in Algorithm 3.
@@
$\hat{\alpha}_{g},\hat{\beta}_{g},\hat{\delta}_{cg},\hat{\gamma}_{cg}$ Cluster ComBat parameters and $k$ trained K-means model
@@
     Estimate local $\hat{\delta}_{ig},\hat{\gamma}_{ig}$ using $\text{EmpricalBayes}(Z_{ijg})$
@@
     Estimate $\delta_{cg}^{*},\gamma_{cg}^{*}$ by taking average of all $\hat{\delta}_{ig},\hat{\gamma}_{ig}$ for all sites $i$ belong to cluster $c$.
@@
$y_{ijg}$ new client’s data, $X_{ij}$ biological covariate of new client, trained cluster-wise harmonization parameters $\hat{\alpha}_{g},\hat{\beta}_{g},\hat{\delta}_{cg},\hat{\gamma}_{cg}$, and trained K-means $k$
@@
Predict cluster index $c=k(\hat{\alpha}_{ig},\hat{\beta}_{ig},\hat{\gamma}_{ig})$
@@
return $y_{ijg}^{*}=\frac{\hat{\sigma}_{g}}{\delta_{cg}^{*}}(Z_{ijg}-\gamma_{cg}^{*})+\hat{\alpha}_{g}+X_{ij}\hat{\beta}_{g}$
@@ 4.1. Synthetic data generation
We follow the data generation procedure in (Silva et al., 2023) and use the graphical model in Figure 1. We replace the site effects with the cluster effects $\gamma_{cg}$. Specifically, the value $y_{ijg}$ with feature index $g$, site index $i$, and data point index $j$ is considered as $y_{ijg}\sim\mathcal{N}(\alpha_{g}+X_{ij}\beta_{g}+\gamma_{cg},\delta_{cg}^{2}\sigma_{g}^{2})$. Note that feature values with index $g$ that are from different sites but in the same cluster will be affected by the same cluster effects $\gamma_{cg}$ and $\delta_{cg}$. The ground truth of the harmonized feature for $y_{ijg}$ (expected feature value after harmonization) is $\alpha_{g}+X_{ij}\beta_{g}$. Besides, we induce the task binary label information into the biological covariate $X_{ij}$. For instance, $X_{ij}\sim\mathcal{N}(0.5,0.5)$ and $X_{ij}\sim\mathcal{N}(-0.5,0.5)$ imply positive and negative labels, respectively, so that the ground truth data is linearly separated. To visualize the site pattern, cluster pattern, and label pattern in the synthetic data, Principal Components Analysis (PCA) is employed to reduce the dimension of data to 2 (Maćkiewicz and Ratajczak, 1993). Figure 2(a) and Figure 2(b) are examples of visualizing the raw (unharmonized) data, which show the patterns of site, cluster, and downstream task labels.
@@
Figure 2. Synthetic Data: site pattern and label pattern of raw data.
@@ 4.2. Synthetic data experiment
We first verify our motivation that if the feature values of sites exhibit cluster patterns in the feature space, then locally estimated feature-wise parameters will also exhibit the same cluster pattern in the parameter space. We generated synthetic data points for 9 sites within 3 clusters for cluster visualization (with data configuration that the number of sites, sample per site, feature, sites per cluster, and biological covariate are 9, 10, 20, 3 and 5 respectively). Specifically, sites 1, 2, and 3 are in the same cluster, sites 4, 5, and 6 are in the same cluster, and sites 7, 8, and 9 are in the same cluster. We used PCA to reduce the dimension to 2 and visualize data points of all sites in the feature space as well as the locally trained parameters for each site in the parameter space. We use colored circles to show the cluster pattern in both feature space and parameter space. As demonstrated in Figure 3, sites within the same cluster in the feature space (as shown in Figure 3(a)) can also be clustered into the same cluster in the parameter space (as shown in Figure 3(b)). This indicates that cluster patterns in the feature space can be retained in the parameter space, which does verify our motivation.
@@
Then, we verify the efficacy of our algorithm over ComBat in both centralized and distributed settings using synthetic data. We generate five synthetic data sets, which follow the graphical model Figure 1 with different parameter configurations. The detailed generation configurations are summarized in Table 1. We assess the performance of Cluster ComBat harmonization and original ComBat algorithms on the synthetic data with aforementioned conditions over two tasks: ground-truth data, i.e., $\alpha_{g}+X_{ij}\beta_{g}$, reconstruction task and ground-truth label classification tasks. Specifically, the Root Mean Square Error (RMSE) between the ground-truth data and the harmonized test data is proposed as the performance measure of the reconstruction task. Also, task accuracy is naturally selected as the performance measure of the downstream classification task. Because the original ComBat and Distributed ComBat cannot harmonize data from unseen sites, we will retrain harmonization parameters whenever they have data from a testing site. Meanwhile, our proposed Cluster ComBat and Distributed Cluster ComBat can harmonize testing data without the need for retraining the ComBat algorithm. In the experiment, we divide synthetic data into 70% for training and 30% for testing for each task and report the mean and variance of performance measures over 30 random seeds. The results are summarized in Table 2. The results show that Cluster ComBat and Distributed Cluster ComBat outperform ComBat and Distributed ComBat in both tasks over various data conditions.
@@
Without harmonization 14.35±0.27 35.65±12.15 22.07±1.07 31.53±9.87 31.67±3.02 96.97±0.01 97.58±0.00 98.42±0.00 98.59±0.00 98.85±0.00
ComBat[a] 6.53±0.03 14.49±1.73 7.40±0.08 11.74±0.90 9.16±0.23 96.92±0.02 90.08±0.14 98.57±0.00 98.65±0.00 99.00±0.00
Cluster ComBat 6.43±0.05 14.38±2.75 7.29±0.11 11.70±1.22 9.06±0.31 97.03±0.01 97.93±0.00 98.59±0.00 98.84±0.00 98.94±0.00
@@
Distributed ComBat[a] 6.60±0.03 14.54±1.58 7.42±0.07 11.75±0.84 9.19±0.21 95.50±0.04 90.65±0.27 97.49±0.01 98.03±0.01 97.34±0.03
Distributed Cluster ComBat 6.44±0.05 14.37±2.62 7.28±0.11 11.69±1.17 9.04±0.31 97.22±0.02 97.68±0.01 98.68±0.00 98.77±0.00 98.94±0.00

### v3

@@ Abstract.
Independent and identically distributed (i.i.d.) data is essential to many data analysis and modeling techniques. In the medical domain, collecting data from multiple sites or institutions is a common strategy that guarantees sufficient clinical diversity, determined by the decentralized nature of medical data. However, data from various sites are easily biased by the local environment or facilities, thereby violating the i.i.d. rule. A common strategy is to harmonize the site bias while retaining important biological information. The ComBat is among the most popular harmonization approaches and has recently been extended to handle distributed sites. However, when faced with situations involving newly joined sites in training or evaluating data from unknown/unseen sites, ComBat lacks compatibility and requires retraining with data from all the sites. The retraining leads to significant computational and logistic overhead that is usually prohibitive. In this work, we develop a novel Cluster ComBat harmonization algorithm, which leverages cluster patterns of the data in different sites and greatly advances the usability of ComBat harmonization. We use extensive simulation and real medical imaging data from ADNI to demonstrate the superiority of the proposed approach. Our codes are provided in https://github.com/illidanlab/distributed-cluster-harmonization.
@@
copyright: acmlicensedjournalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671590isbn: 979-8-4007-0490-1/24/08ccs: Applied computing Imagingccs: Computing methodologies Distributed algorithmsccs: Computing methodologies Machine learning
@@ 1. Introduction
The recent advances in machine learning approaches have greatly advanced biomedical data analysis. In brain imaging analysis, for example, Magnetic Resonance Imaging (MRI) has been used for the detection and disease progression of many diseases, such as Mild Cognitive Impairment (MCI) (Wang et al., 2018; Zhou et al., 2011, 2013, 2012), Parkinson’s disease (Garcia Santa Cruz et al., 2023), and Brain Tumor Detection (Abdusalomov et al., 2023). However, one critical challenge with brain imaging is that the brain imaging is sensitive to scanner or protocol effect (Zhan et al., 2014; Pfaehler et al., 2019), also commonly referred to as site effect or batch effect, leading to the fact that brain imaging from multiple sites is not independent and identically distributed (i.i.d.). The bias in non-i.i.d. data will cause unstable prediction performance and poor generalization performance to unseen data (Yang et al., 2021). Consequently, developing an algorithm that can eliminate these types of bias ensures consistent and reliable outcomes in the deployment of machine learning models within the medical imaging domain.
@@
ComBat (Johnson et al., 2006) is a well-known harmonization technique and has been shown to be helpful in mitigating the site effect of neuroimaging data introduced by multiple sites sampling (Orlhac et al., 2021). Despite its utility, one of the central ideas is that ComBat debias the site effect independently according to the (local) site data, which induces one critical limitation, its inability to evaluate site effects coming from unseen or unknown sites without undergoing a retraining process. The requirement of retraining is hindered by substantial computational costs when it comes to real-world deployment, especially when dealing with large datasets from multiple sites. This limitation underscores the need for a more efficient and broadly applicable approach in mitigating the site effects of medical data, especially neuroimaging data.
@@
Furthermore, a centralized setting for site effect harmonization introduces extra concerns. For instance, sharing data directly among multiple sites to apply ComBat harmonization poses challenges to the security of confidential data and the protection of patient privacy. Direct training on all the data is often impractical in the medical domain. This underscores the need to develop harmonization algorithms in a decentralized manner that can effectively harmonize without gathering data from all sites while maintaining competitive performance in the centralized setting. Particularly, the Distributed ComBat (Chen et al., 2022), a distributed version of the ComBat harmonization algorithm, has demonstrated its harmonization capability meanwhile obeying decentralized manners. Nevertheless, Distributed ComBat suffers from the same limitation as the original ComBat, i.e., it cannot estimate site effects from unseen sites without retraining.
@@
Because a significant part of site effects in medical data are ultimately rooted in medical instruments, e.g., MRI scanners from different manufacturers and configurations, the bias underneath the sites may not be independent and may exhibit clustering structures. In this paper, we proposed the Cluster ComBat method, an extension of the original ComBat algorithms that leverages the cluster patterns of site effects. This approach enables the estimation of site effects from unknown sites without necessitating a retraining process. Furthermore, we also developed a distributed version Cluster ComBat and demonstrated its efficacy in harmonizing data obeying decentralized manners. Our empirical findings show that Cluster ComBat in both centralized and decentralized settings outperform their respective counterparts on both synthetic and real-world neuroimaging datasets.
@@ 2. Related Works
Distributed Learning. Preserving the privacy of users’ information is a crucial issue that needs to be considered as an important aspect to evaluate in a learning algorithm  (Li et al., 2021; Fukasawa et al., 2004; Xie et al., 2018; Hong et al., 2022, 2021). Especially in a healthcare setting, where the health data is sensitive, we need to design a distributed learning approach that avoids leaking any private information from hospitals’ data  (Xie et al., 2017; Welten et al., 2022; Wirth et al., 2021; Haque et al., 2023). For example, for brain imaging, federated learning, a distributed machine learning algorithm, has proven to be effective in analyzing neuroimaging for cognitive detection tasks while still protecting patients’ information  (Sandhu et al., 2023; Dipro et al., 2022). Moreover, for health records data, to avoid sharing raw data between sites, the use of first-order and second-order gradients of the likelihood function of sites has been shown to be sufficient for achieving high accuracy in classification tasks (Duan et al., 2019). In addition, the distributed version of generalized linear mixed models (GLMM) can also achieve nearly identical results in analyzing electronic health records data as in a centralized setting  (Yan et al., 2023). Relating to harmonization methods in a decentralized setting, Distributed ComBat (Chen et al., 2022) and Federated Learning ComBat (Silva et al., 2023) have been developed to harmonize neuroimaging without the need for sharing information between hospitals.
@@ 3.1. Preliminary: ComBat Harmonization
(3) $\gamma_{ig}^{*}=\frac{N_{i}\overline{\tau}_{i}^{2}\hat{\gamma}_{ig}+{\delta_{ig}^{*}}^{2}\overline{\gamma}_{i}}{N_{i}\overline{\tau}_{i}^{2}+{\delta_{ig}^{*}}^{2}},\quad\quad{\delta_{ig}^{*}}^{2}=\frac{\overline{\theta}_{i}+\frac{1}{2}\sum_{j}(Z_{ijg}-\gamma_{ig}^{*})^{2}}{\frac{1}{2}N_{i}+\overline{\lambda}_{i}-1},$
where $\overline{\tau}_{i}^{2}$, $\overline{\gamma}_{i}$, $\overline{\theta_{i}}$, $\overline{\lambda_{i}}$ are computed through the method of moments.
@@ 3.2. Cluster ComBat
Using cluster-wise shared harmonization parameters, we can generalize knowledge from previous $N$ sites to the new unseen site once we know which cluster each data point from this site belongs to. Additionally, the estimation process of harmonization parameters $\gamma_{cg}$ and $\delta_{cg}$ can benefit from multiple sites’ data points within the same cluster, considering the sample number for estimating each parameter group is enlarged. And we name this algorithm Cluster ComBat.
@@
$\hat{\alpha}_{g},\hat{\beta}_{g},\delta_{cg}^{*},\gamma_{cg}^{*}$ harmonization parameters and $k$ trained K-means model
@@
$y_{ijg}$ unseen site’s data, $X_{ij}$ biological covariate of new client, previously estimated parameters $\hat{\alpha}_{g},\hat{\beta}_{g},\delta_{cg}^{*},\gamma_{cg}^{*}$, and trained K-means $k$
@@ 3.3. Distributed Cluster ComBat
The Distributed Cluster ComBat has the following steps: i) each site estimates feature-wise parameters $\hat{\alpha}_{ig}$, $\hat{\beta}_{ig}$, and $\hat{\gamma}_{ig}$ locally at the same time, and sends parameters to the central server;  ii) the central server performs K-means clustering based on $\hat{\alpha}_{ig}$, $\hat{\beta}_{ig}$, and $\hat{\gamma}_{ig}$ for $i\in[M]$;  iii) the central server aggregates $\{\hat{\alpha}_{ig}\}_{i\in M}$, $\{\hat{\beta}_{ig}\}_{i\in M}$ and $\{\hat{\gamma}_{ig}\}_{i\in M}$ to estimates the global feature-wise parameters $\hat{\alpha}_{g}$, $\hat{\beta}_{g}$ and $\hat{\gamma}_{ig}$, and then sends back to each site;  iv) each site standardized the local data using global feature-wise parameters, then locally estimates $\hat{\delta}_{ig}$ and $\hat{\gamma}_{ig}$;  v) each site sends locally estimated harmonization parameters to the server;  vi) server aggregates harmonization parameters within each cluster to estimate the cluster-wise ones, then sends back to each site; the aggregation procedure precisely follows the procedure outlined in Figure 1 of the original Distributed ComBat algorithm paper (Chen et al., 2022);  vii) each site performs local harmonization based on cluster-wise harmonization parameters. The procedure is summarized in Algorithm 3.
@@
$\hat{\alpha}_{g},\hat{\beta}_{g},\delta_{cg}^{*},\gamma_{cg}^{*}$ Cluster ComBat parameters and $k$ trained K-means model
@@
     Estimate local $\delta_{ig}^{*},\gamma_{ig}^{*}$ using $\text{EmpricalBayes}(Z_{ijg})$ via Equation 3
@@
     Estimate $\delta_{cg}^{*},\gamma_{cg}^{*}$ by taking average of all $\delta_{ig}^{*},\gamma_{ig}^{*}$ for all sites $i$ belong to cluster $c$.
@@
$y_{ijg}$ new client’s data, $X_{ij}$ biological covariate of new client, trained cluster-wise harmonization parameters $\hat{\alpha}_{g},\hat{\beta}_{g},\delta_{cg}^{*},\gamma_{cg}^{*}$, and trained K-means $k$
@@
Get cluster index $\tilde{c}=k(\hat{\alpha}_{ig},\hat{\beta}_{ig},\hat{\gamma}_{ig})$
@@ 4.1. Synthetic data generation
We follow the data generation procedure in (Silva et al., 2023) and use the graphical model in Figure 1. We replace the site effects with the cluster effects $\gamma_{cg}$. Specifically, the value $y_{ijg}$ with feature index $g$, site index $i$, and data point index $j$ is considered as $y_{ijg}\sim\mathcal{N}(\alpha_{g}+X_{ij}\beta_{g}+\gamma_{cg},\delta_{cg}^{2}\sigma_{g}^{2})$. Note that feature values with index $g$ that are from different sites but in the same cluster will be affected by the same cluster effects $\gamma_{cg}$ and $\delta_{cg}$. The ground truth of the harmonized feature for $y_{ijg}$ (expected feature value after harmonization) is $\alpha_{g}+X_{ij}\beta_{g}$. Besides, we induce the task binary label information into the biological covariate $X_{ij}$. For instance, $X_{ij}\sim\mathcal{N}(0.5,0.5)$ and $X_{ij}\sim\mathcal{N}(-0.5,0.5)$ imply positive and negative labels, respectively, so that the ground truth data is linearly separated. To visualize the site pattern, cluster pattern, and label pattern in the synthetic data, Principal Components Analysis (PCA) is employed to reduce the dimension of the data to 2 (Maćkiewicz and Ratajczak, 1993). Figure 2(a) and Figure 2(b) are examples of visualizing the raw (unharmonized) data, which show the patterns of site, cluster, and downstream task labels.
@@
Figure 2. Synthetic Data: site pattern and label pattern of the raw data.
@@ 4.2. Synthetic data experiment
We first verify our motivation that if the feature values of sites exhibit cluster patterns in the feature space, then locally estimated feature-wise parameters will also exhibit the same cluster pattern in the parameter space. We generated synthetic data points for 9 sites within 3 clusters for cluster visualization (with data configuration that the number of sites, sample per site, feature, sites per cluster, and biological covariate are 9, 10, 20, 3, and 5 respectively). Specifically, sites 1, 2, and 3 are in the same cluster, sites 4, 5, and 6 are in the same cluster, and sites 7, 8, and 9 are in the same cluster. We used PCA to reduce the dimension to 2 and visualize data points of all sites in the feature space as well as the locally trained parameters for each site in the parameter space. We use colored circles to show the cluster pattern in both feature space and parameter space. As demonstrated in Figure 3, sites within the same cluster in the feature space (as shown in Figure 3(a)) can also be clustered into the same cluster in the parameter space (as shown in Figure 3(b)). This indicates that cluster patterns in the feature space can be retained in the parameter space, which verifies our motivation.
@@
Then, we verify the efficacy of our algorithm over ComBat in both centralized and distributed settings using synthetic data. We generate five synthetic data sets, which follow the graphical model Figure 1 with different parameter configurations. The detailed generation configurations are summarized in Table 1. We assess the performance of Cluster ComBat harmonization and original ComBat algorithms on the synthetic data with aforementioned conditions over two tasks: ground-truth data, i.e., $\alpha_{g}+X_{ij}\beta_{g}$, reconstruction task and ground-truth label classification tasks. Specifically, the Root Mean Square Error (RMSE) between the ground-truth data and the harmonized test data is proposed as the performance measure of the reconstruction task. Also, task accuracy is naturally selected as the performance measure of the downstream classification task. Because the original ComBat and Distributed ComBat cannot harmonize data from unseen sites, we will retrain harmonization parameters whenever they have data from a testing site. Meanwhile, our proposed Cluster ComBat and Distributed Cluster ComBat can harmonize testing data without the need for retraining the ComBat algorithm. In the experiment, we divided the synthetic data into 70% for training and 30% for testing for each task and reported the mean and variance of performance measures over 30 random seeds. The results are summarized in Table 2. The results show that Cluster ComBat and Distributed Cluster ComBat outperform ComBat and Distributed ComBat in both tasks over various data conditions.
@@
Without harmonization 14.35±0.28 35.65±12.15 22.07±1.07 31.53±9.87 31.67±3.03 96.97±0.02 97.57±0.01 98.42±0.00 98.59±0.00 98.85±0.00
ComBat[a] 6.53±0.03 14.49±1.73 7.40±0.08 11.74±0.90 9.16±0.23 96.92±0.02 90.10±0.14 98.57±0.00 98.65±0.00 99.00±0.00
Cluster ComBat 6.43±0.05 14.38±2.75 7.29±0.12 11.70±1.22 9.06±0.32 97.03±0.01 97.93±0.00 98.59±0.00 98.84±0.00 98.94±0.00
@@
Distributed ComBat[a] 6.60±0.03 14.54±1.58 7.42±0.07 11.75±0.84 9.19±0.21 95.53±0.04 90.65±0.27 97.47±0.01 98.02±0.01 97.35±0.03
Distributed Cluster ComBat 6.44±0.05 14.37±2.62 7.28±0.11 11.69±1.17 9.04±0.31 97.22±0.02 97.70±0.01 98.68±0.00 98.77±0.00 98.93±0.00